A radiator contains 5 quarts of fluid, 10% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 50% antifreeze?
Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39617)
Answer by richwmiller(17219)

0.1*5-0.1x+1x=0.5(5)
0.5+0.9x=2.5
0.9x=2.5-0.5
0.9x=2
x=2.22222222
check
0.1*2.77777778+1*2.22222222=0.5(5)
0.27777778+2.22222222=2.5
2.5=2.5
ok
coderep